# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v1

@@ 7 Limitations
We highlight several limitations of our work that we left for future explorations. First, in order to compute path integration, we sum the rotation angles along the temporal dimension, meaning that our formalism only applies to causal transformers. Further work is needed to generalize our method to non-causal problems such as image processing, but the same methods used to adapt SSMs to vision zhu2024visionmambaefficientvisual could be used. Second, we did not try to scale our method to larger model sizes and larger datasets, but since MapFormers require little extra compute compared to baseline transformers, computation time should not be a bottleneck. Third, our results only confirm the superiority of EM models on recall tasks compared to their WM counterparts, but further experiments should compare these two models on more complex tasks that involve reasoning. Finally, in our models, for numerical efficiency, we only use block diagonal matrices, without orthogonal matrix $M$ used in eq.2, since orthogonal matrices do not change the value of a scalar product and should have no impact on the attention. However, real neurons cannot have a negative firing rate, and because of that, position $p_{t}$ must always be positive, which is impossible with a block diagonal rotation matrix. Matrix $M$ mixes frequencies together to make the code non negative dorrell2023actionableneuralrepresentationsgrid, which creates a fundamental difference between our models and biological systems. Non-negativity and bounded activity can also result in disentangled representations whittington2023disentanglementbiologicalconstraintstheory, which could help build even more robust and interpretable cognitive maps in the future.
@@ 10.1 MAmPa: Mamba with skew-symmetric block-diagonal matrices
As explained in sec.3.5, one only needs to modify the recurrent matrix $A:=S=-S^{\top}$ of selective SSMs to be block-diagonal skew-symmetric in order to learn spatial cognitive maps. However, the reason why people chose diagonal matrices in Mamba is to dramatically reduce the inference speed of these models, and make them as close as possible to parallelized. In Mamba, an efficient implementation of the parallel scan is used, leveraging the fact that $A$ is diagonal and can therefore be treated as a vector. Even then, the author reduce the internal state to $N=16$, which has a dramatic impact on the ability to learn cognitive maps, as demonstrated in fig. 4. In Mamba2, the authors go even further by choosing $A:=a\in\mathbb{R}$ to be a scalar, which allows them to broadcast it easily in the attention matrix and increase the state size to $N=64$ (eq. 16).
However, as we have demonstrated in our experiments, diagonal matrices are not expressive enough to learn spatial cognitive maps, nor any other higher dimensional map. We implemented a parallel scan on square matrices to test if MAmPa, an extended version of Mamba with $2\times 2$ block-diagonal matrices can learn spatial maps like our theory predicts. Because the parallel scan on bigger matrices dramatically increases the computational load, we were forced to keep the inner state small ($N=16$) and only trained models on sequences of size $l=16$. We report the results in tab. 3, where OOD dense and OOD sparse use sequences of size $8$ and $32$ respectively.
@@
This confirms that our formalism extends to SSMs, but also shows that what has allowed the scaling of selective SSMs – diagonal matrices to reduce the computational load – is what prevents them from learning cognitive maps, leaving only transformers as scalable implementations of structure learning models.

### v2

@@ 7 Limitations
We highlight several limitations of our work that we left for future explorations. First, in order to compute path integration, we sum the rotation angles along the temporal dimension, meaning that our formalism only applies to causal transformers. Further work is needed to generalize our method to non-causal problems such as visual processing, but the same methods used to adapt SSMs to vision zhu2024visionmambaefficientvisual could be used. Second, we did not compare MapFormers to PoPE gopalakrishnan2025decouplingwhatwherepolar, but since it relies on a mechanism similar to CoPE, we expect it to behave similarly. Third, we did not try to scale our method to larger model sizes and larger datasets, but since MapFormers require little extra compute compared to baseline transformers, computation time should not be a bottleneck. Fourth, our results only confirm the superiority of EM models on recall tasks compared to their WM counterparts, but further experiments should compare these two models on more complex tasks. Finally, in our models, for numerical efficiency, we only use block diagonal matrices, without orthogonal matrix $M$ used in eq. 2. However, real neurons cannot have a negative firing rate, and because of that, position $p_{t}$ must always be positive, which is impossible with a block diagonal rotation matrix. Matrix $M$ mixes frequencies together to make the code non negative dorrell2023actionableneuralrepresentationsgrid, which creates a fundamental difference between our models and biological systems. Non-negativity and bounded activity can also result in disentangled representations whittington2023disentanglementbiologicalconstraintstheory, which could help build even more robust and interpretable cognitive maps in the future.
@@ 10.1 MAmPa: Mamba with skew-symmetric block-diagonal matrices
As explained in sec.3.4, one only needs to modify the recurrent matrix $A:=S=-S^{\top}$ of selective SSMs to be block-diagonal skew-symmetric in order to learn spatial cognitive maps. However, the reason why people chose diagonal matrices in Mamba is to dramatically reduce the inference speed of these models, and make them as close as possible to parallelized. In Mamba, an efficient implementation of the parallel scan is used, leveraging the fact that $A$ is diagonal and can therefore be treated as a vector. In Mamba2, the authors go even further by choosing $A:=a\in\mathbb{R}$ to be a scalar, which allows them to broadcast it easily in the attention matrix and increase the model’s dimension (eq. 16), showing that recent attempts at scaling SSMs mainly rely on decreasing the selective weights expressivity, and therefore learning of cognitive maps at scale.
However, as we have demonstrated in our experiments, diagonal matrices are not expressive enough to learn spatial cognitive maps, nor any other higher dimensional map. We implemented a parallel scan on square matrices to test if MAmPa, an extended version of Mamba with $2\times 2$ block-diagonal matrices can learn spatial maps like our theory predicts. Because the parallel scan on square matrices dramatically increases the computational load, we were forced to keep the inner state small ($N=16$) and only trained models on sequences of size $l=16$ since we did not take the time to implement a custom parallel scan using the efficient implementation of matrix-vector product for rotations used in RoPE (eq. 18), but we expect it to be highly beneficial to lean cognitive maps at scale with Mamba models. We report the results in tab. 3, where OOD dense and OOD sparse use sequences of size $8$ and $32$ respectively.
@@
This confirms that our formalism extends to SSMs, but also shows that what has allowed the scaling of selective SSMs – diagonal matrices to reduce the computational load – is what prevents them from learning cognitive maps.